# FlexDiT: Dynamic Token Density Control for Diffusion Transformer

### v1

@@ 3.2 FlexDiT
Given $T$ sampling steps, we propose a sample-specific approach to dynamically adjust the pruning rate $r$, thereby controlling the number of sparse tokens across sampling steps. We define a range for $r$, such that $r\in[r_{\text{min}},r_{\text{max}}]$. Observing that generation quality is highly relative with the later denoising stages (in Section 4.4), we hold $r$ constant at $r_{\text{min}}$ for the first $T/4$ sampling steps. Subsequently, we adjust the pruning rate $r$ linearly based on the current sampling step $t_{i}$. The specific formula for $r$ is provided as follows:
@@
During training, we sample step $t_{i}$ and compute the corresponding $r$ according to Eq. 6. However, the tokens of input should be same to train the model in batch, and the randomness of sampling from $T$ also need to be maintained. To solve this contradiction, we modify the linear function of Eq. 6 when $T/4\leq t_{i}<T$ to piecewise function. Normally, the model is trained by multiple GPUs. We request the $t_{i}$ sampled in a specific piece in each GPU. Therefore, in each iteration, it can achieve both uniformly random sampling and training in batch.
3.4 Initialization and fine-tuning.
Our method fine-tunes pre-trained DiT models to improve efficiency. The processes of generating sparse tokens and restoring dense tokens utilize off-the-shelf transformers, avoiding the need for additional networks. During fine-tuning, the parameters of transformers in DiT-based model are loaded into the corresponding transformers in our FlexDiT, with two exceptions: (1) queries and keys are absent in Poolingformers, hence related parameters in pre-trained model are not required; (2) the weights $W_{1}$ and $W_{2}$ in Eq. 4 are initialized by full zeros and identity matrix, receptively.
@@
(b) Images generated by FlexDiT (-43% FLOPs).
Figure 4: Generating images from DiT-XL and FlexDiT-XL with the same random seed.
@@
Table 2: FVD scores of FlexDiT-XL in class-conditional video generation task on four mainstream video generation datasets. We do not find the official reference sets of these 4 dataset when we compute FVD scores, hence we build the reference sets by ourselves and use the official checkpoint to re-compute the FVD scores of Latte-XL in this table.
@@ 4 Experiments
FlexDiT is applied in three representative DiT-based models, DiT [40], Latte [33], and PixArt-$\alpha$ [8] for class-conditional image generation and video generation, and text-to-image generation, respectively.
@@ 4.1 Class-conditional image generation
We conduct our experiments on ImageNet-1k [11] at resolutions of $256\times 256$ and $512\times 512$, following the protocol established in DiT. For DiT-XL, the model consists of 2, 24, and 2 transformer layers in the bottom, middle, and top segments, respectively. The middle segment includes 4 sparse-dense token modules, which comprise 1 sparse token generation transformer, 3 sparse transformers, 1 dense token recovery transformer, and 1 dense transformer. For DiT-B, the bottom, middle, and top segments contain 1, 10, and 1 transformer layers, respectively. The middle segment consists of 2 sparse-dense token modules, including 1 sparse token generation transformer, 2 sparse transformers, 1 dense token recovery transformer, and 1 dense transformer.
As poolingformers and sparse transformers can interfere with the function of position embedding, we reintroduce sine-cosine position embedding at each stage of dense token generation. For DiT-XL, we use the checkpoint from the official DiT repository, while for DiT-B, we utilize the checkpoint provided by [39]. All training settings and hyperparameters follow those in the original DiT paper. The strength of classifier-free guidance [18] is set to 1.5 for evaluation and 4.0 for visualization.
Following prior works, we sample 50,000 images to compute the Fréchet Inception Distance (FID) [17] using the ADM TensorFlow evaluation suite [12], along with the Inception Score (IS) [47], sFID [37], and Precision-Recall metrics [23]. Throughput is evaluated with a batch size of 128 on an Nvidia A100 GPU.
@@
The results of our approach on the class-conditional image generation task using ImageNet are presented in Table 1. We evaluate our method on two model sizes, DiT-B and DiT-XL, and at two resolutions, $256\times 256$ and $512\times 512$. At pruning rates $r\in[0.61,0.86]$, FlexDiT-XL achieves a 43% reduction in FLOPs and an 87% improvement in inference speed, with only a 0.11 increase in FID score. This indicates significant redundancy within DiT, as using only about 25% tokens in specific layers maintains similar performance levels. By increasing the token count, we observe slight improvements over baseline performance with 145% improvement in throughput. The impact of our approach is even more pronounced at a resolution of $512\times 512$. With a higher pruning rate than the $256\times 256$ resolution, our method yields a superior performance-efficiency trade-off. By pruning over 90% of the tokens, we achieve a 55% reduction in FLOPs and a 175% increase in throughput, with only a 0.09 increase in FID score. When reducing the pruning rate further, our results slightly surpass the baseline, delivering a 145% throughput improvement.
From $256\times 256$ to $512\times 512$, the FLOPs increase by only 4.4 times, whereas the speed declines by a factor of 6.3, demonstrating that the primary speed bottleneck in DiT structure is the number of tokens. Our method effectively reduces token count, resulting in significant gains, particularly for high-resolution content generation. For instance, we achieve a 55% reduction in FLOPs at $512\times 512$ while enhancing speed by 175%. Compared to other methods, our real-world speed improvements far exceed those achieved with similar reductions in FLOPs.
@@
To further validate the effectiveness of our method, we visualize some samples generated from FlexDiT-XL at a $256\times 256$ resolution with a pruning rate $r\in[0.61,0.86]$ (-43% FLOPs) and compare them with DiT-XL in Figure 4. Each cell in the same position corresponds to images generated from the same random seed and class. Since our model is fine-tuned from the pre-trained DiT, the overall styles and structures of the two subfigures are similar. The images generated by FlexDiT still retain rich high-frequency details. Furthermore, FlexDiT images demonstrate more precise structure, as evidenced by the accurate depiction of the “golden retriever’s” nose and the “macaw’s” eyes, which appear misplacement or missing in the images generated by DiT-XL. We provide samples at a resolution of $512\times 512$ in Supplementary [sec:vis]..
@@ 4.2 Class-conditional video generation
We conduct experiments at a resolution of $256\times 256$ on four public datasets: FaceForensics [45], SkyTimelapse [60], UCF101 [54], and Taichi-HD [50]. Latte comprises two distinct types of transformer blocks: spatial transformers, which focus on capturing spatial information, and temporal transformers, which capture temporal information. To accommodate this separation of spatial and temporal feature extraction, we adjust our model’s schedule within the sparse-dense token module. Specifically, we employ two transformers to prune spatial tokens and temporal tokens separately. The sparse tokens are then processed by two sparse transformers, and finally, two transformers are used to recover the temporal and spatial tokens. Dense transformers are discarded in this setup. All other model configurations are the same as in DiT, and we follow the original training settings and hyperparameters used for Latte.
To evaluate performance, we sample 2,048 videos, each consisting of 16 frames, and measure the Fréchet Video Distance (FVD) [57]. Throughput is measured using a batch size of 2 clips on an Nvidia A100 GPU.
@@ 4.3 Text-to-image generation
Experiment setting
We further assess the effectiveness of our method on text-to-image generation, which presents a greater challenge compared to class-conditional image generation. We adopt PixArt-$\alpha$ [8], a text-to-image generation model built based on DiT as the base model. Our model is initialized using the official PixArt-$\alpha$ checkpoint pre-trained on SAM dataset [22] containing 10M images. We further fine-tune it with our method on a subset of SAM dataset including 1M images. In PixArt-$\alpha$, the transformer architecture includes two types of attention layers: a self-attention layer and a cross-attention layer, which integrates textual information. We apply our method to the self-attention layers to either reduce or recover tokens.
For evaluation, we randomly select text prompts from SAM dataset and adopt 100-step IDDPM solver [38] to sample 30,000 images, with the FID score serving as the evaluation metric. Throughput is evaluated with a batch size of 128 on an Nvidia A100 GPU.
Main results
As shown in Table 3, our FlexDiT achieves an FID score comparable to the original PixArT-$\alpha$ with significantly accelerating the generation, showing that our method is effective for text-to-image generation task and shows a strong general-purpose ability in various generation fields.
@@
(b) Images generated by FlexDiT
Figure 5: Generating images from PixArt-$\alpha$ (a) and FlexDiT (b). The captions in the same positions are same. The captions are presented in Supplementary 7.
@@ 4.4 Ablation study
All the following ablation experiments of FlexDiT-XL are conducted on ImageNet at a $256\times 256$ resolution.
@@
The alternation between sparse and dense tokens is a key factor contributing to the success of our method. Sparse tokens efficiently capture global structures, while dense tokens capture detailed information. In FlexDiT, we adopt four SDTMs by default. Table 4 shows the effect of using different numbers of SDTMs. To isolate the impact of SDTM count, we maintain the same FLOPs across models by adjusting the number of sparse and dense transformers while varying only the number of SDTMs. In configurations with only one SDTM, the model is modified into a U-shaped architecture, allocating 6 transformers to the bottom segment, 16 to a single SDTM, and 6 to the top segment. Note that in this configuration, standard transformers are used in the bottom layers instead of poolingformers prevent any performance degradation that arises from the use of poolingformers. The results indicate that fewer SDTMs reduce interactions between global and local information, resulting in weaker performance. Furthermore, when FlexDiT is reduced to a U-Net structure, training stability is compromised, resulting in training collapse. We do not experiment with more than four SDTMs, as increasing SDTM count beyond four offers diminishing returns in FLOP reduction. Using four SDTMs achieves an optimal balance between performance and efficiency.
No. of SDTMs 1 2 3 4
FID $\downarrow$ NAN 3.86 2.51 2.38
FLOPs 68.74 67.00 69.97 68.05
Table 4: Performance evaluation on different numbers of SDTMs. “NAN” indicates loss nan during training.
No. of Poolingformers 0 1 2 3
No. of trans in top seg. 4 3 2 1
FID $\downarrow$ NAN 2.38 2.38 2.56
FLOPs 69.56 68.80 68.05 67.29
Table 5: Performance evaluation on different numbers of poolingformers. “NAN” indicates loss nan during training.
Model 250-DDPM 25-DDIM 5-RFlow
DiT-XL 2.27 2.89 43.40
Ours 2.38 3.31 43.69
Table 6: Combination our method with efficient samplers. The numbers represent FID scores.
FLOPs (G) No. of Tokens Dynamic FID $\downarrow$
$\sim 68$ $8\times 8$ 2.48
$6\times 6\sim 10\times 10$ 2.38
$\sim 61$ $6\times 6$ 2.63
$4\times 4\sim 8\times 8$ 2.50
Table 7: Comparison of FID scores with and without timestep-wise pruning rate strategy.
@@
Combination with efficient samplers.
Our FlexDiT is a model compression method that can be seamlessly integrated with efficient samplers, such as DDIM [51] and Rectified Flow (RFlow) [28]. Note that the RFlow variant of DiT was trained by us via 2, [^0] K iterations. As shown in Table 6, when our method is combined with the 25-step DDIM and 5-step Rectified Flow samplers, it achieves approximately 18.7×18.7$\times$18.7 × and 93.4×93.4$\times$93.4 × improvements in inference speed, respectively, compared to the standard 250-step DDPM. Notably, our method does not introduce a significant performance gap relative to the baseline, demonstrating that it can be effectively combined with efficient samplers.
@@ 5 Limitations
6 Conclusion
In this paper, we introduced FlexDiT, a novel method for enhancing the efficiency of DiT by leveraging token sparsity. Through a strategically layered approach that incorporates SDTM and timestep-wise pruning rate, FlexDiT significantly reduces computational complexity while maintaining high generation quality. By decoupling global and local feature extraction with sparse and dense tokens, FlexDiT achieves efficient utilization of resources in both image and video generation tasks. Our experimental results demonstrate substantial reductions in FLOPs with minimal performance degradation, validating FlexDiT’s effectiveness across multiple datasets. Moreover, FlexDiT complements existing methods aimed at optimizing sampling steps, enabling further improvements when used in tandem with enhanced ODE solvers. This work presents a step forward in developing scalable, efficient diffusion models suitable for high-resolution content generation, opening up new possibilities for practical applications of DiT-based architectures.
Supplementary Material
7 Captions in Figure 5

### v2

@@ 3.2 Timestep-wise pruning rate
Given $T$ sampling steps, we propose a sample-specific approach to dynamically adjust the pruning rate $r$, thereby controlling the number of sparse tokens across the sampling steps. We define a range for $r$, such that $r\in[r_{\text{min}},r_{\text{max}}]$. Observing that generation quality is highly relative with the later denoising stages (in Section 4.4), we hold $r$ constant at $r_{\text{min}}$ for the first $T/4$ sampling steps. Subsequently, we adjust the pruning rate $r$ linearly based on the current sampling step $t_{i}$. The specific formula for $r$ is provided as follows:
@@
During training, we sample step $t_{i}$ and compute the corresponding $r$ according to Eq. 5. However, the input tokens should be the same to train the model in batch, and the randomness of the sampling from $T$ should also be maintained. To solve this contradiction, we modify the linear function of Eq. 5 when $T/4\leq t_{i}<T$ to a piecewise function. Normally, the model is trained by multiple GPUs. We request $t_{i}$ sampled in a specific piece in each GPU. Therefore, in each iteration, it can achieve both uniformly random sampling and batch training.
3.3 Initialization and fine-tuning
Our method fine-tunes pre-trained DiT models to improve efficiency. The processes of generating sparse tokens and restoring dense tokens utilize off-the-shelf transformers, avoiding the need for additional networks. During fine-tuning, the parameters of transformers in DiT-based model are loaded into the corresponding transformers in our SparseDiT, with two exceptions: (1) queries and keys are absent in Poolingformers, hence related parameters in pre-trained model are not required; (2) the weights $W_{1}$ and $W_{2}$ in Eq. 3 are initialized by full zeros and identity matrix, receptively.
@@ 3.3 Initialization and fine-tuning
(b) Images generated by SparseDiT (-43% FLOPs).
Figure 4: Generating images from DiT-XL and SparseDiT-XL with the same random seed.
@@
Table 2: FVD scores of SparseDiT-XL in class-conditional video generation task on four mainstream video generation datasets. We do not find the official reference sets of these 4 dataset when we compute FVD scores, hence we build the reference sets by ourselves and use the official checkpoint to re-compute the FVD scores of Latte-XL in this table.
@@ 4 Experiments
SparseDiT is applied in three representative DiT-based models, DiT peebles2023scalable , Latte ma2024latte , and PixArt-$\alpha$ chen2023pixartalpha for class-conditional image generation, class-conditional video generation, and text-to-image generation, respectively.
Fine-tuning overhead.
All training settings and hyperparameters follow their respective papers. Fine-tuning requires approximately 6% of the time needed for training from scratch, e.g., 400K iterations for DiT-XL fine-tuning.
@@ 4.1 Class-conditional image generation
We conduct our experiments on ImageNet-1k deng2009imagenet at resolutions of $256\times 256$ and $512\times 512$, following the protocol established in DiT. For DiT-XL, the model consists of 2, 24, and 2 transformers in the bottom, middle, and top segments, respectively. The middle segment includes 4 sparse-dense token modules, which comprise 1 sparse token generation transformer, 3 sparse transformers, 1 dense token recovery transformer, and 1 dense transformer. For DiT-B, the bottom, middle, and top segments contain 1, 10, and 1 transformer layers, respectively. The middle segment consists of 2 sparse-dense token modules, including 1 sparse token generation transformer, 2 sparse transformers, 1 dense token recovery transformer, and 1 dense transformer. As poolingformers and sparse transformers can interfere with the function of position embedding, we reintroduce sine-cosine position embedding at each stage of dense token generation. For DiT-XL, we use the checkpoint from the official DiT repository, while for DiT-B, we utilize the checkpoint provided by pan2024t . Following prior works, we sample 50,000 images to compute the Fréchet Inception Distance (FID) heusel2017gans using the ADM TensorFlow evaluation suite dhariwal2021diffusion , along with the Inception Score (IS) salimans2016improved , sFID nash2021generating , and Precision-Recall metrics kynkaanniemi2019improved . Classifier-free guidance ho2022classifier (CFG) is set to 1.5 for evaluation and 4.0 for visualization. Throughput is evaluated with a batch size of 128 on an Nvidia A100 GPU.
@@
The results of our approach on the class-conditional image generation task using ImageNet are presented in Table 1. We evaluate our method on two model sizes, DiT-B and DiT-XL, and at two resolutions, $256\times 256$ and $512\times 512$. At pruning rates $r\in[0.61,0.86]$, SparseDiT-XL achieves a 43% reduction in FLOPs and an 87% improvement in inference speed, with only a 0.11 increase in FID score. This indicates significant redundancy within DiT, as using only about 25% tokens in specific layers maintains similar performance levels. By increasing the token count, we observe slight improvements over baseline performance with 145% improvement in throughput. The impact of our approach is even more pronounced at a resolution of $512\times 512$. With a higher pruning rate than the $256\times 256$ resolution, our method yields a superior performance-efficiency trade-off. By pruning over 90% of the tokens, we achieve a 55% reduction in FLOPs and a 175% increase in throughput, with only a 0.09 increase in FID score. When reducing the pruning rate further, our results slightly surpass the baseline, delivering a 145% throughput improvement. ToMeSD bolya2023tomesd is the first approach to reduce tokens. However, it applies token merging and recovery at every layer in a generic manner, resulting in a significant performance drop. For example, with a 0.1 merging rate on DiT-XL, ToMeSD achieves an FID score of 14.74, which is substantially worse than our result. A detailed comparison between other methods and our method is provided in Appendix A.1.
From $256\times 256$ to $512\times 512$, FLOPs increase by only 4.4 times, whereas the speed declines by a factor of 6.3, demonstrating that the primary speed bottleneck in DiT architecture is the number of tokens. Our method effectively reduces token count, resulting in significant gains, particularly for high-resolution content generation. For example, we achieve a 55% reduction in FLOPs at $512\times 512$ while enhancing speed by 175%. Compared to other methods, our real-world speed improvements far exceed those achieved with similar reductions in FLOPs.
@@
To further validate the effectiveness of our method, we visualize some samples generated from SparseDiT-XL at a $256\times 256$ resolution with a pruning rate $r\in[0.61,0.86]$ (-43% FLOPs) and compare them with DiT-XL in Figure 4. Each cell in the same position corresponds to images generated from the same random seed and class. Since our model is fine-tuned from the pre-trained DiT, the overall styles and structures of the two subfigures are similar. The images generated by SparseDiT still retain rich high-frequency details. Furthermore, SparseDiT images demonstrate more precise structure, as evidenced by the accurate depiction of the “golden retriever’s” nose and the “macaw’s” eyes, which appear misplacement or missing in the images generated by DiT-XL. We provide samples at a resolution of $512\times 512$ in Appendix.
@@ 4.2 Class-conditional video generation
We conduct experiments at a resolution of $256\times 256$ on four public datasets: FaceForensics rossler2018faceforensics , SkyTimelapse xiong2018learning , UCF101 soomro2012dataset , and Taichi-HD siarohin2019first . To evaluate performance, we sample 2,048 videos, each consisting of 16 frames, and measure the Fréchet Video Distance (FVD) unterthiner2018towards . Throughput is measured using a batch size of 2 clips on an Nvidia A100 GPU. More experimental settings are shown in Appendix A.3.
@@ 4.3 Text-to-image generation
Experiment setting.
We further assess the effectiveness of our method on text-to-image generation, which presents a greater challenge compared to class-conditional image generation. We adopt PixArt-$\alpha$ chen2023pixartalpha , a text-to-image generation model built based on DiT as the base model. Our model is initialized using the official PixArt-$\alpha$ checkpoint pre-trained on SAM dataset kirillov2023segment containing 10M images. We further fine-tune it with our method on a subset of SAM dataset including 1M images. In PixArt-$\alpha$, the transformer architecture includes two types of attention layers: a self-attention layer and a cross-attention layer, which integrates textual information. We apply our method to the self-attention layers to either reduce or recover tokens.
For evaluation, we randomly select text prompts from SAM dataset and adopt 100-step IDDPM solver nichol2021improved to sample 30,000 images, with the FID score serving as the evaluation metric. Throughput is evaluated with a batch size of 128 on an Nvidia A100 GPU.
Main results.
As shown in Table 3, our SparseDiT achieves an FID score comparable to the original PixArT-$\alpha$ with significantly accelerating the generation, showing that our method is effective for text-to-image generation task.
@@
(b) Images generated by SparseDiT
Figure 5: Generating images from PixArt-$\alpha$ (a) and SparseDiT (b). The captions in the same positions are identical. They are provided in Appendix A.4.
@@ 4.4 Ablation study
All the following ablation experiments of SparseDiT-XL are conducted on ImageNet at a $256\times 256$ resolution.
@@
The alternation between sparse and dense tokens is a key factor contributing to the success of our method. Sparse tokens capture global structures, while dense tokens capture detailed information. In SparseDiT, we adopt four SDTMs by default. Table 9 shows the effect of using different numbers of SDTMs. To isolate the impact of SDTM count, we maintain the same FLOPs across models by adjusting the number of sparse and dense transformers while varying only the number of SDTMs. In configurations with only one SDTM, the model is modified into a U-shaped architecture. The results indicate that fewer SDTMs reduce interactions between global and local information, resulting in weaker performance. Furthermore, reducing SparseDiT to a U-Net structure compromises training stability, leading to a collapse. We do not experiment with more than four SDTMs, as increasing SDTM count beyond four offers diminishing returns in FLOP reduction.
[…]
Combination with efficient samplers.
Our SparseDiT is a model compression method that can be seamlessly integrated with efficient samplers, such as DDIM song2020denoising and Rectified Flow (RFlow) liu2022flow . Note that the RFlow variant of DiT was trained by us via 2, [^0] K iterations. As shown in Table 9, when our method is combined with the 25-step DDIM and 5-step Rectified Flow samplers, it achieves approximately $18.7\times$ and $93.4\times$ improvements in inference speed, respectively, compared to the standard 250-step DDPM. Notably, our method does not introduce a significant performance gap relative to the baseline, demonstrating that it can be effectively combined with efficient samplers.
@@ 5 Conclusion
In this paper, we introduce SparseDiT, a novel approach to enhancing the efficiency of DiT by leveraging token sparsity. By incorporating SDTM and a timestep-wise pruning rate in a strategically layered manner, SparseDiT significantly reduces computational complexity while preserving high-generation quality. Experimental results show substantial FLOP reductions with minimal performance degradation, demonstrating SparseDiT’s effectiveness across multiple DiT-based models and datasets. This work marks a step forward in developing scalable and efficient diffusion models for high-resolution content generation, expanding the practical applications of DiT-based architectures.
Appendix A Appendix
A.1 Comparison with other methods
Comparison with ToMeSD.
Recent studies moon2023early ; zhao2024dynamic have evaluated ToMeSD bolya2023tomesd on DiT models and reported significant performance drops, which are much lower than those achieved by our method. Our approach is fundamentally different from ToMeSD. While ToMeSD reduces and recovers tokens generically at every layer, our method specifically analyzes the local-global relationships unique to DiT models. Based on these insights, we strategically reduce and recover tokens. We have evaluated ToMeSD on DiT and compared it with our method, as shown in Table 16.
Comparison with DyDiT.
DyDiT zhao2024dynamic represents the current state-of-the-art approach for enhancing the efficiency of Diffusion Transformers (DiTs). In contrast to our method, which emphasizes token reduction, DyDiT implements pruning across tokens, attention heads, and channels. A comparative analysis of our method against DyDiT, as detailed in Table 17 using ImageNet 512×512 images, illustrates a substantial advantage in computational efficiency for our approach. Specifically, our method achieves a 46% reduction in FLOPs, markedly surpassing DyDiT’s 29% reduction, coupled with a 145% increase in inference speed, significantly outperforming the 31% improvement achieved by DyDiT. The only trade-off is a minor increase in the Fréchet Inception Distance (FID), where our method’s FID is 0.08 higher than that of DyDiT. Nevertheless, this difference in FID is nearly imperceptible to human observers in practical applications.
Comparison with TokenCache.
We compare our method with TokenCache lou2024token in Table 17. We adopt the standard DiT baseline 2.27, while TokenCache applies a baseline 2.24. TokenCache achieves a 32% speed gain with a slight FID increase. Our approach attains an 87% speed gain with a smaller FID increase, demonstrating superior efficiency and performance.
Comparison with Ditfastattn.
We compare our method with Ditfastattn yuan2025ditfastattn in Table 17. We adopt the standard DiT baseline 3.04, while Ditfastattn applies a baseline 3.16. Ditfastattn shows a substantial FID increase for speed gain, while our method results in better speed and less FID.
A.2 Training form scratch
We train our SparseDiT from scratch, reaching 400K training iterations. We compare our result with original DiT paper in Table 4. The results (cfg=1.0) are in Table 18.
A.3 More experimental settings in Sparse-Latte
Latte comprises two distinct types of transformer blocks: spatial transformers, which focus on capturing spatial information, and temporal transformers, which capture temporal information. To accommodate this separation of spatial and temporal feature extraction, we adjust our model’s schedule within the sparse-dense token module. Specifically, we employ two transformers to prune spatial tokens and temporal tokens separately. The sparse tokens are then processed by two sparse transformers, and finally two transformers are used to recover the temporal and spatial tokens. Dense transformers are discarded in this setup. All other model configurations are the same as in DiT, and we follow the original training settings and hyperparameters used for Latte.
A.4 Captions in Figure 5